You want to purchase a few pen and pencils from Ms.Barad. If 8 pens and 7 pencils cost $3.37 while 5 pens and 11 pencils cost $3.10. How much does each pen and each pencil cost?
Answer by mallikarjuna1st(3)

1)8x + 7y = 3.37
2)5x + 11y = 3.10
multiply 1) by 5 and 2) by 8
40x + 35y = 16.85
40x + 88y = 24.80
----------------------
0x - 53y = -7.95
therefore y = 0.15
Sub y = 0.15 in 1)
8x = 3.37 - 1.05
x = .29
Answer is Pen = 0.29$ and Pencil = 0.15$